what is 5 X - 15 y equals 30 converted to slope intercept form with work so I can learn from it

Answer by amarjeeth123(569)

5x-15y=30
15y=5x-30
y=1/3x-2
y=(1/3)x+(-2)
The slope is 1/3 and the intercept is -2.